# Locally Optimal Fixed-Budget Best Arm Identification in Two-Armed Gaussian Bandits with Unknown Variances

### 20 Dec 2023 v1

@@ 1 Introduction
Formally, we consider the following adaptive experiment with two arms and Gaussian rewards. There are two arms $1$ and $2$, and an arm $a\in\{1,2\}$ has an $\mathbb{R}$-valued Gaussian reward $Y_{a}\sim\mathcal{N}(\mu_{a},\sigma_{a}^{2})$ with the mean $\mu_{a}\in\mathbb{R}$ and the variance $\sigma_{a}^{2}>0$. Let $P=P(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=(\mathcal{N}(\mu_{1},\sigma^{2}_{1}),\mathcal{N}(\mu_{2},\sigma^{2}_{2}))$ be a pair of distributions that generate $(Y_{1},Y_{2})$, where $\mathcal{N}(\mu,\sigma^{2})$ is a Gaussian distribution with a mean $\mu$ and a variance $\sigma^{2}$. Let $\mathcal{P}^{\mathrm{G}}=\{P(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}):\mu_{1}\neq\mu_{2}\}$ be a set of the distributions, which is referred to as the Gaussian bandit models. For an instance $P$, the best arm $a^{\star}(P)\in\{1,2\}$ is defined as $a^{\star}(P)=\operatorname*{arg\,max}_{a\in\{1,2\}}\mu_{a}$, which is assumed to exist uniquely.
In the adaptive experiment, we consider a strategy to identify the best arm. A fixed budget $T$ is given. For each round $t\in[T]:=\{1,2,\dots,T\}$, let $(Y_{1,t},Y_{2,t})$ be an independent and identically distributed (i.i.d.) copy of $(Y_{1},Y_{2})$. At each round $t$, we draw arm $A_{t}\in\{1,2\}$ and observe a reward $Y_{t}=\sum_{a\in\{1,2\}}\mathbbm{1}[A_{t}=a]Y_{a,t}$. At the end of the experiment (after round $T$), we recommend an estimated best arm $\widehat{a}_{T}\in\{1,2\}$. During an experiment, we follow a strategy that determines which arm to draw and which arm to recommend as the best arm. The performance of strategies is evaluated by a minimal probability of misidentification $\mathbb{P}_{P}(\widehat{a}_{T}\neq a^{\star}(P))$, where $\mathbb{P}_{P}$ is the probability law under $P$.
@@
Based on the situation above, our interest is in strategies for identifying misidentification probabilities in the adaptive experimental setting described above. Specifically, we need a strategy such that an upper bound on its misidentification probability is aligns with the lower bound proposed in Kaufmann et al. (2016). Further, this strategy must be valid when the variance is unknown.
@@
The small-gap regime considers a situation $\mu_{1}-\mu_{2}\to 0$ as $T\to\infty$. Although this additional setting slightly simplifies the problem with BAI, the problem is still sufficiently complicated since the small gap makes it difficult to identify the best arm. This setting has been utilized in BAI with fixed confidence, such as the analysis of lil’UCB (Jamieson et al., 2014). In statistical test, such an evaluation framework is known as the local Bahadur efficiency (Bahadur, 1960; Wieand, 1976; Akritas & Kourouklis, 1988; He & Shao, 1996). . From a technical perspective, the small-gap regime is a situation where we can ignore the estimation error of the variances compared to the difficulty of identifying the best arm. Since the error of the estimation of the variance is relatively negligible in the small-gap setting, we can show that the misidentification probability of the NA-AIPW strategy matches the lower bound.
@@
We note some relation to the existing studies on the BAI problem. Because we restrict our analysis to the small-gap setting, our result does not contradict the analysis on another lower bound of the misidentification probability by Ariu et al. (2021) and Degenne (2023). We discuss this problem in Section 4.
@@
This study is organized as follows. First, in Section 2, we review the lower bound of Kaufmann et al. (2016). Then, in Section 3, we propose our NA-AIPW strategy. In Section 4, we show that the misidentification probability of the strategy asymptotically corresponds to the lower bound by Kaufmann et al. (2016) under the small-gap setting. We show the proof in Section 5, where we also provide a novel concentration inequality based on the Chernoff bound. In Section 6, we discuss related work and remaining problems, which includes an extension of our small-gap setting to a setting with multi-armed bandits and non-Gaussian rewards.
Notation. Let $\mathcal{F}_{t}$ be the sigma-algebra generated by all observations up to round $t$. We define a truncation operator: for a variable $v\in\mathbb{R}$ and a constant $c\geq 1$, $\mathrm{thre}(v;c):=\max\{\min\{v,c\},1/c\}$.
@@ 2 Lower Bound of Probability of Misidentification
As a preparation, we introduce a lower bound for the probability of misidentification in BAI with a fixed budget. We call a strategy is consistent, if for any $P\in\mathcal{P}^{\mathrm{G}}$, $\mathbb{P}_{P}(\widehat{a}_{T}\neq a^{\star}(P))\to 0$ as $T\to\infty$. To evaluate the performance of strategies for each $P\in\mathcal{P}^{\mathrm{G}}$, we focus on the following metric for $\mathbb{P}_{P}(\widehat{a}_{T}\neq a^{\star}(P))$ used in many studies, such as Kaufmann et al. (2016):
@@
For two-armed Gaussian bandits, Kaufmann et al. (2016) presents the following lower bounds. Let $P^{*}\in\mathcal{P}^{\mathrm{G}}$ be a distribution that generates the rewards $\{(Y_{1,t},Y_{2,t})\}_{t\in[T]}$.
@@
For each $P^{*}=(\mathcal{N}(\mu^{*}_{1},\sigma^{2}_{1}),\mathcal{N}(\mu^{*}_{2},\sigma^{2}_{2})\in\mathcal{P}^{\mathrm{G}}$ and $\Delta=\mu^{*}_{1}-\mu^{*}_{2}$, any consistent strategy satisfies
$\limsup_{T\to\infty}-\frac{1}{T}\log\mathbb{P}_{P^{*}}(\widehat{a}_{T}\neq a^{\star}(P^{*}))\leq\frac{\Delta^{2}}{2\big{(}\sqrt{\sigma^{2}_{1}}+\sqrt{\sigma^{2}_{2}}\big{)}^{2}}.$
@@
It has been discussed to find a strategy in which the upper bound of its probability of misidentification coincides with this lower bound in Proposition 2.1. Although Kaufmann et al. (2016) develops a strategy that satisfies the requirement, it needs to sample arms with some probability depending on the known variances $(\sigma_{1}^{2},\sigma_{2}^{2})$. To the best of our knowledge, if the variances are unknown and need to be estimated during adaptive experiments, no one has found the desired strategy. For this open problem, also see Kaufmann (2020), Ariu et al. (2021), Qin (2022), and Degenne (2023).
@@ 3.1 Target Allocation Ratio
$w^{*}_{1}=\frac{{\sigma_{1}}}{{\sigma^{2}_{1}}+{\sigma_{2}}},\mbox{~{}~{}and~{}~{}}w^{*}_{2}=1-w^{*}_{1}.$
@@ 3.2 Sampling Rule with Neyman Allocation (NA)
$\widetilde{\mu}_{a,t}=\frac{1}{\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]}\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]Y_{a,s}.$
Also, we define a second moment estimator $\widetilde{\zeta}_{a,t}=({\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]})^{-1}\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]Y^{2}_{a,s}$, and a root of variance estimator $\widetilde{\sigma}_{a,t}=\{\widetilde{\zeta}_{a,t}-\left(\widetilde{\mu}_{a,t}\right)^{2}\}^{1/2}$. Then, we define the estinator $\widehat{\sigma}_{a,t}=\mathrm{thre}(\widetilde{\sigma}_{a,t};C_{\sigma^{2}}^{1/2})$ with some predetermined constant $C_{\sigma^{2}}>0$. Also, we define the estimator $\widehat{w}_{1,t}$ and $\widehat{w}_{2,t}$ as
$\widehat{w}_{1,t}=\frac{{\widehat{\sigma}_{1,t}}}{{\widehat{\sigma}_{1,t}}+{\widehat{\sigma}_{2,t}}},\quad\mathrm{and}\quad\widehat{w}_{2,t}=1-\widehat{w}_{1,t}.$ (1)
At the end of the round $t\geq 3$, we sample the arm $A_{t}$; we generate $\gamma_{t}$ from the uniform distribution on $[0,1]$ and set
$A_{t}=\begin{cases}1&\mbox{if~{}}\gamma_{t}\leq\widehat{w}_{1,t}\\ 2&\mbox{otherwise}.\end{cases}$
We note that it is possible to increase the number of rounds for initialization, while we present that only the first two rounds are used for initialization. The more rounds for initialization have the role of stabilizing the sampling rule in practice. This idea is similar to the forced-sampling (Garivier & Kaufmann, 2016).
@@ 3.3 Recommendation Rule with the Augmented Inverse Probability Weighting (AIPW) Estimator
We present our recommendation rule. In the recommendation phase after round $T$, we estimate $\mu^{*}_{a}$ for each $a\in\{1,2\}$ and recommend an arm with the bigger estimated expected reward. With a truncated version of the estimated expected reward $\widehat{\mu}_{a,t}=\mathrm{thre}(\widetilde{\mu}_{a,t},C_{\mu})$ with some predetermined constant $C_{\mu}>0$, we define the augmented inverse probability weighting (AIPW) estimator of $\mu^{*}_{a}$ for each $a\in\{1,2\}$ as
$\widehat{\mu}^{\mathrm{AIPW}}_{a,T}=\frac{1}{T}\sum^{T}_{t=1}\psi_{a,t},\qquad\mathrm{where}\ \psi_{a,t}=\frac{\mathbbm{1}[A_{t}=a]\big{(}Y_{a,t}-\widehat{\mu}_{a,t}\big{)}}{\widehat{w}_{a,t}}+\widehat{\mu}_{a,t}.$ (2)
@@
$\widehat{a}_{T}=\begin{cases}&1\quad\mathrm{if}\quad\widehat{\mu}^{\mathrm{AIPW}}_{1,T}\geq\widehat{\mu}^{\mathrm{AIPW}}_{2,T},\\ &2\quad\mathrm{otherwise}.\end{cases}$ (3)
We adopt the AIPW estimator for our strategy because it has several advantages. First, the AIPW estimator has the property of semiparametric efficiency, which indicates that it has the smallest asymptotic variance among a certain class (Hahn, 1998). The property is necessary to prove that the strategy using the AIPW estimator is optimal, which means the misidentification probability is small enough to achieve its lower bound. The second reason is more technical; the AIPW estimator simplifies a proof for theoretical analysis. Specifically, we can decomposed an error by the AIPW estimator into a sum of random variables with martingale properties, making it suitable for analysis using the central limit theorem. This property is unique to the AIPW estimator, but not to naive estimators such as an empirical average. Details will be given in Section 5.
@@
     Sample $\gamma_{t}$ from the uniform distribution on $[0,1]$.
     $A_{t}=1$ if $\gamma_{t}\leq\widehat{w}_{1,t}$; $A_{t}=2$ if $\gamma_{t}>\widehat{w}_{1,t}$.
@@ 4 Misidentification Probability and Asymptotic Optimality
For each $P^{*}\in\mathcal{P}^{\mathrm{G}}$, the following holds as $\Delta\to 0$:
$\liminf_{T\to\infty}-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)\geq\frac{\Delta^{2}}{2(\sigma_{1}+\sigma_{2})^{2}}+O\left(\Delta^{3}\right).$
Note that the lower bound of $-\log\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)$ implies the upper bound of $\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)$. This theorem implies us to evaluate the probability of misidentification up to the constant term, even when it is exponentially small, as $\Delta\to 0$.
@@
To show Theorem 4.1, we derive the upper bound of $\mathbb{P}_{P^{*}}(\widehat{\mu}^{\mathrm{AIPW}}_{a^{\star}(P^{*}),T}\leq\widehat{\mu}^{\mathrm{AIPW}}_{b,T})$ for $b\in\{1,2\}\backslash\{a^{\star}(P^{*})\}$, which is equivalent to $\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)$. Without loss of generality, we assume that $a^{\star}(P^{*})=1$ and $b=2$. Let us define $V=\frac{\sigma^{2}_{1}}{w^{*}_{1}}+\frac{\sigma^{2}_{2}}{w^{*}_{2}}=\left(\sigma_{1}+\sigma_{2}\right)^{2}$ and
$\Psi_{t}$ $=\frac{\psi_{1,t}-\psi_{2,t}-\Delta}{\sqrt{V}}.$
@@
Step 1: Sequence $\{\Psi_{t}\}^{T}_{t=1}$ is an MDS
@@
$\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\sum^{T}_{t=1}\Psi_{t}\right)\right]=\mathbb{E}_{P^{*}}\left[\prod^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]\right]=\mathbb{E}_{P^{*}}\left[\exp\left(\sum^{T}_{t=1}\log\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]\right)\right].$ (4)
@@
$\log\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]=\frac{\lambda^{2}}{2}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]+O\left(\lambda^{3}\right).$ (5)
Here, $\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]=1+\sum^{\infty}_{k=1}\lambda^{k}\mathbb{E}_{P^{*}}\left[\Psi^{k}_{t}/k!|\mathcal{F}_{t-1}\right]$. Because $\mathbb{E}_{P^{*}}\left[\Psi^{k}_{t}/k!|\mathcal{F}_{t-1}\right]$ is bounded by a constant that is independent of $T$ for all $k\geq 1$,
$\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]=1+\sum^{2}_{k=1}\lambda^{k}\mathbb{E}_{P^{*}}\left[\Psi^{k}_{t}/k!|\mathcal{F}_{t-1}\right]+O\left(\lambda^{3}\right).$
Note that the Taylor series expansion of $\log(1+z)$ around $z=0$ is given as $\log(1+z)=z-z^{2}/2+z^{3}/3-\cdots$. Therefore, we have

### 17 Mar 2024 v2

@@ 1 Introduction
Formally, we consider the following adaptive experiment with two arms and Gaussian rewards. There are two arms $1$ and $2$, and an arm $a\in\{1,2\}$ has an $\mathbb{R}$-valued Gaussian reward $Y_{a}\sim\mathcal{N}(\mu_{a},\sigma_{a}^{2})$ with the mean $\mu_{a}\in[-C_{\mu},C_{\mu}]$ and the variance $\sigma_{a}^{2}\in[C_{\sigma^{2}},1/C_{\sigma^{2}}]$ for some universal constants $C_{\mu},C_{\sigma^{2}}>0$. We assume that $C_{\mu}$ and $C_{\sigma^{2}}$ are known to us for a technical purpose, and it is enough to set $C_{\mu}$ as a sufficiently large value and $C_{\sigma^{2}}$ as a sufficiently small value. Given fixed $(\sigma^{2}_{1},\sigma^{2}_{2})$, let
$\mathcal{P}^{\mathrm{G}}\coloneqq\mathcal{P}^{\mathrm{G}}_{(\sigma^{2}_{1},\sigma^{2}_{2})}\coloneqq\big{\{}P=(\mathcal{N}(\mu_{1},\sigma^{2}_{1}),\mathcal{N}(\mu_{2},\sigma^{2}_{2})):\mu_{1},\mu_{2}\in(-\infty,+\infty),\ \ \mu_{1}\neq\mu_{2}\big{\}}$
be a set of distributions generating the data, which is referred to as the Gaussian bandit models, where $P\in\mathcal{P}$ is a pair of distributions that generate $(Y_{1},Y_{2})$, and $\mathcal{N}(\mu,\sigma^{2})$ is a Gaussian distribution with a mean $\mu$ and a variance $\sigma^{2}$. For an instance $P$, the best arm $a^{\star}(P)\in\{1,2\}$ is defined as $a^{\star}(P)=\operatorname*{arg\,max}_{a\in\{1,2\}}\mu_{a}$, which is assumed to exist uniquely.
In the adaptive experiment, we consider a strategy to identify the best arm. A fixed budget $T$ is given. For each round $t\in[T]\coloneqq\{1,2,\dots,T\}$, let $(Y_{1,t},Y_{2,t})$ be an independent and identically distributed (i.i.d.) copy of $(Y_{1},Y_{2})$. At each round $t$, we draw arm $A_{t}\in\{1,2\}$ and observe a reward $Y_{t}=\sum_{a\in\{1,2\}}\mathbbm{1}[A_{t}=a]Y_{a,t}$. At the end of the experiment (after round $T$), we recommend an estimated best arm $\widehat{a}_{T}\in\{1,2\}$. During an experiment, we follow a strategy that determines which arm to draw and which arm to recommend as the best arm. The performance of strategies is evaluated by a minimal probability of misidentification $\mathbb{P}_{P}(\widehat{a}_{T}\neq a^{\star}(P))$, where $\mathbb{P}_{P}$ is the probability law under $P$.
@@
Based on the situation above, our interest is in strategies for identifying misidentification probabilities in the adaptive experimental setting described above. Specifically, we need a strategy such that an upper bound on its misidentification probability aligns with the lower bound proposed in Kaufmann et al. (2016). Further, this strategy must be valid when the variance is unknown.
@@
The small-gap regime considers a situation $\mu_{1}-\mu_{2}\to 0$ as $T\to\infty$. Although this additional setting slightly simplifies the problem with BAI, the problem is still sufficiently complicated since the small gap makes it difficult to identify the best arm. This setting has been utilized in BAI with fixed confidence, such as the analysis of lil’UCB (Jamieson et al., 2014). In the realm of statistical testing, such an evaluation framework is known as the local Bahadur efficiency (Bahadur, 1960; Wieand, 1976; Akritas & Kourouklis, 1988; He & Shao, 1996). From a technical perspective, the small-gap regime is a situation where we can ignore the estimation error of the variances compared to the difficulty of identifying the best arm. Since the error of the estimation of the variance is relatively negligible in the small-gap setting, we can show that the misidentification probability of the NA-AIPW strategy matches the lower bound.
@@
This study is organized as follows. First, in Section 2, we review the lower bound of Kaufmann et al. (2016). Then, in Section 3, we propose our NA-AIPW strategy. In Section 4, we show that the misidentification probability of the strategy asymptotically corresponds to the lower bound by Kaufmann et al. (2016) under the small-gap setting. We show the proof in Section 5, where we also provide a novel concentration inequality based on the Chernoff bound. In Section 6, we discuss the difficulty in this problem. In Section 7, we introduce related work and remaining problems, which includes an extension of our small-gap setting to a setting with multi-armed bandits and non-Gaussian rewards.
Notation. Let $\mathcal{F}_{t}$ be the sigma-algebra generated by all observations up to round $t$. We define a truncation operator: for a variable $v\in\mathbb{R}$ and a constant $c\geq 1$, $\mathrm{thre}(v;c_{1},c_{2})\coloneqq\min\{\max\{v,c_{1}\},c_{2}\}$.
@@ 2 Lower Bound of Probability of Misidentification
As a preparation, we introduce a lower bound for the probability of misidentification in BAI with a fixed budget. We call a strategy is consistent, if for any $P\in\mathcal{P}^{\mathrm{G}}$, $\mathbb{P}_{P}(\widehat{a}_{T}\neq a^{\star}(P))\to 0$ as $T\to\infty$. To evaluate the performance of strategies For any $P\in\mathcal{P}^{\mathrm{G}}$, we focus on the following metric for $\mathbb{P}_{P}(\widehat{a}_{T}\neq a^{\star}(P))$ used in many studies, such as Kaufmann et al. (2016):
@@
For two-armed Gaussian bandits, Kaufmann et al. (2016) presents the following lower bounds.
@@
For any $P^{*}\in\mathcal{P}^{\mathrm{G}}$ and $\Delta=\mu^{*}_{1}-\mu^{*}_{2}$ with known constants $C_{\mu},C_{\sigma^{2}}>0$ independent of $T$, if $\{(Y_{1,t},Y_{2,t})\}_{t\in[T]}$ is generated from $P^{*}$, any consistent strategy satisfies
$\limsup_{T\to\infty}-\frac{1}{T}\log\mathbb{P}_{P^{*}}(\widehat{a}_{T}\neq a^{\star}(P^{*}))\leq\frac{\Delta^{2}}{2\big{(}\sigma_{1}+\sigma_{2}\big{)}^{2}}.$
Note that we can remove the condition that we know constants $C_{\mu},C_{\sigma^{2}}>0$ independent of $T$ such that $\mu_{a}\in[-C_{\mu},C_{\mu}]$ and $\sigma^{2}_{1},\sigma^{2}_{2}>C_{\sigma^{2}}$ hold for deriving the lower bound. However, it is required to implement our strategy and to derive an upper bound. For the conditions of the lower bound to align with those of the upper bound, we add the conditions in this proposition.
@@
It has been discussed to find a strategy in which the upper bound of its probability of misidentification coincides with this lower bound in Proposition 2.1. Although Kaufmann et al. (2016) develops a strategy that satisfies the requirement, it needs to sample arms with some probability depending on the known variances $(\sigma_{1}^{2},\sigma_{2}^{2})$. To the best of our knowledge, if the variances are unknown and need to be estimated during adaptive experiments, no one has found the desired strategy.
@@ 3.1 Target Allocation Ratio
$w^{*}_{1}\coloneqq\frac{{\sigma_{1}}}{{\sigma_{1}}+{\sigma_{2}}},\mbox{~{}~{}and~{}~{}}w^{*}_{2}\coloneqq 1-w^{*}_{1}.$
@@ 3.2 Sampling Rule with Neyman Allocation (NA)
$\widetilde{\mu}_{a,t}\coloneqq\frac{1}{\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]}\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]Y_{a,s}.$
Also, we define a second moment estimator $\widetilde{\zeta}_{a,t}\coloneqq({\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]})^{-1}\sum^{t-1}_{s=1}\mathbbm{1}[A_{s}=a]Y^{2}_{a,s}$, and a root of variance estimator $\widetilde{\sigma}_{a,t}=\{\widetilde{\zeta}_{a,t}-\left(\widetilde{\mu}_{a,t}\right)^{2}\}^{1/2}$. Then, we define the estimator $\widehat{\sigma}_{a,t}=\mathrm{thre}(\widetilde{\sigma}_{a,t};\underline{C}_{\sigma^{2}}^{1/2},1/\overline{C}_{\sigma^{2}})$ with the constant $C_{\sigma^{2}}>0$, defined in Section 1. Note that this truncation is introduced for a technical purpose to draw each arm infinitely many times as $T\to\infty$ and avoid the estimators of $\mu^{*}_{1}$ and $\mu^{*}_{2}$, defined below, diverging to infinity. We just use a sufficiently small value for $C_{\sigma^{2}}>0$. Also, we define the estimator $\widehat{w}_{1,t}$ and $\widehat{w}_{2,t}$ as
$\widehat{w}_{1,t}\coloneqq\frac{{\widehat{\sigma}_{1,t}}}{{\widehat{\sigma}_{1,t}}+{\widehat{\sigma}_{2,t}}},\quad\mathrm{and}\quad\widehat{w}_{2,t}\coloneqq 1-\widehat{w}_{1,t}.$ (1)
In each round $t\geq 3$, we draw arm $A_{t}=1$ with probability $\widehat{w}_{1,t}$ and $A_{t}=2$ with probability $\widehat{w}_{2,t}$.
We note the possibility of increasing the number of initialization rounds, although our strategy utilizes only the first two rounds for this purpose. The additional rounds of initialization serve to stabilize the sampling rule in practical applications, akin to the concept of the forced-sampling (Garivier & Kaufmann, 2016). We can change the number of initialization rounds if the condition $*\widehat{w}_{a,t}\xrightarrow{\mathrm{a.s.}}w^{*}_{a}$ is satisfied as $t\to\infty$ for every $a\in\{1,2\}$, which is crucial for our theoretical analysis. For instance, instead of using $\widehat{w}_{1,t}$ directly, an alternative arm-drawing probability could be defined as $\widetilde{w}_{1,t}=\alpha_{t}/2+(1-\alpha_{t})\widehat{w}_{1,t}$, assuming $\alpha_{t}\in[0,1]$ and converges to zero as $t$ approaches infinity (here, we define $\widetilde{w}_{2,t}=1-\widetilde{w}_{1,t}$). Moreover, the number of initialization rounds can be made dependent upon the number of arms without impacting the theoretical outcomes.
@@ 3.3 Recommendation Rule with the Augmented Inverse Probability Weighting (AIPW) Estimator
We present our recommendation rule. In the recommendation phase after round $T$, we estimate $\mu^{*}_{a}$ for each $a\in\{1,2\}$ and recommend an arm with the bigger estimated expected reward. With a truncated version of the estimated expected reward $\widehat{\mu}_{a,t}\coloneqq\mathrm{thre}(\widetilde{\mu}_{a,t},-C_{\mu},C_{\mu})$ with some predetermined constant $C_{\mu}>0$, defined in Section 1, we define the augmented inverse probability weighting (AIPW) estimator of $\mu^{*}_{a}$ for each $a\in\{1,2\}$ as
$\widehat{\mu}^{\mathrm{AIPW}}_{a,T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\psi_{a,t},\qquad\mathrm{where}\ \psi_{a,t}\coloneqq\frac{\mathbbm{1}[A_{t}=a]\big{(}Y_{a,t}-\widehat{\mu}_{a,t}\big{)}}{\widehat{w}_{a,t}}+\widehat{\mu}_{a,t}.$ (2)
@@
$\widehat{a}_{T}\coloneqq\begin{cases}&1\quad\mathrm{if}\quad\widehat{\mu}^{\mathrm{AIPW}}_{1,T}\geq\widehat{\mu}^{\mathrm{AIPW}}_{2,T},\\ &2\quad\mathrm{otherwise}.\end{cases}$ (3)
We adopt the AIPW estimator for our strategy because it has several advantages. First, the AIPW estimator has the property of semiparametric efficiency, which indicates that it has the smallest asymptotic variance among a certain class (Hahn, 1998). The property is necessary to prove that the strategy using the AIPW estimator is optimal, which means the misidentification probability is small enough to achieve its lower bound. The second reason is more technical; the AIPW estimator simplifies the theoretical analysis (see Section 6.3). Specifically, we can decompose an error by the AIPW estimator into a sum of random variables with martingale properties, making it suitable for analysis using the central limit theorem. This property is unique to the AIPW estimator but not to naive estimators such as an empirical average. Details will be given in Section 5.
@@
     Draw $A_{t}=1$ with probability $\widehat{w}_{1,t}$ and $A_{t}=2$ with probability $\widehat{w}_{2,t}=1-\widehat{w}_{1,t}$.
@@ 4 Misidentification Probability and Asymptotic Optimality
For any $P^{*}\in\mathcal{P}^{\mathrm{G}}$ with known constants $C_{\mu},C_{\sigma^{2}}>0$ independent of $T$, if $\{(Y_{1,t},Y_{2,t})\}_{t\in[T]}$ is generated from $P^{*}$, the following holds as $\Delta\to 0$:
$\liminf_{T\to\infty}-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)\geq\frac{\Delta^{2}}{2(\sigma_{1}+\sigma_{2})^{2}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-o\left(\Delta^{2}\right)}.$
We note again that $C_{\mu}$ and $C_{\sigma^{2}}$ are introduced for technical purpose. The constant $C_{\mu}$ is introduced to guarantee the boundedness of the estimators, and it is sufficient to use a sufficiently large value for it. The constant $C_{\sigma^{2}}$ is used to draw each arm infinitely many times as $T\to\infty$ and avoid the estimators of the means $\mu^{*}1$ and $\mu^{*}_{2}$ diverging to infinity, and it is sufficient to use a sufficiently small value for $C_{\sigma^{2}}>0$.
Note that the lower bound of $-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)$ implies the upper bound of $\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)$. This theorem implies us to evaluate the probability of misidentification up to the constant term, even when it is exponentially small, as $\Delta\to 0$.
@@
To show Theorem 4.1, we derive the upper bound of $\mathbb{P}_{P^{*}}\left(\widehat{\mu}^{\mathrm{AIPW}}_{a^{\star}(P^{*}),T}\leq\widehat{\mu}^{\mathrm{AIPW}}_{b,T}\right)$ for $b\in\{1,2\}\backslash\{a^{\star}(P^{*})\}$, which is equivalent to $\mathbb{P}_{P^{*}}\left(\widehat{a}^{\mathrm{AIPW}}_{T}\neq a^{\star}(P^{*})\right)$. Without loss of generality, we assume that $a^{\star}(P^{*})=1$ and $b=2$. Let us define $V\coloneqq\frac{\sigma^{2}_{1}}{w^{*}_{1}}+\frac{\sigma^{2}_{2}}{w^{*}_{2}}=\left(\sigma_{1}+\sigma_{2}\right)^{2}$ and
$\Psi_{t}\coloneqq\frac{\psi_{1,t}-\psi_{2,t}-\Delta}{\sqrt{V}}.$
@@
Step 1: Sequence $\{\Psi_{t}\}^{T}_{t=1}$ is a martingale difference sequence (MDS)
@@
$\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\sum^{T}_{t=1}\Psi_{t}\right)\right]=\mathbb{E}_{P^{*}}\left[\prod^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]\right]=\mathbb{E}_{P^{*}}\left[\exp\left(\sum^{T}_{t=1}\log\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]\right)\right].$
@@
$\log\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]=\frac{\lambda^{2}}{2}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]+o\left(\lambda^{2}\right)$ (4)
as $\lambda\to 0$. This is given as follows. Since $\mathbb{E}_{P^{*}}\left[\Psi^{k}_{t}\exp(\lambda\Psi_{t})|\mathcal{F}_{t-1}\right]$ are finite everywhere in an open interval $(0,\infty)$ for $k=1,2,3$, the Taylor expansion yields the following (for the details, see the textbook such as page 75 in Bulmer (1967) and Theorem 5.19 in Apostol (1974)):
$\mathbb{E}_{P^{*}}\left[\exp\left(\lambda\Psi_{t}\right)|\mathcal{F}_{t-1}\right]=1+\sum^{2}_{k=1}\mathbb{E}_{P^{*}}\left[\Psi^{k}_{t}/k!|\mathcal{F}_{t-1}\right]+o\left(\lambda^{2}\right)$
as $\lambda\to 0$. Note that the finiteness of $\mathbb{E}_{P^{*}}\left[\Psi^{k}_{t}\exp(\lambda\Psi_{t})|\mathcal{F}_{t-1}\right]$ comes from the following in $\Psi_{t}$: (i) $Y_{a,t}$ is a Gaussian random variable, (ii) $\widehat{\mu}_{a,t}$ and $\widehat{w}_{a,t}$ are bounded random variables by our truncation, and (iii) the lower bound of $\widehat{w}$ is given by $C_{\sigma^{2}}$. By using the Taylor expansion again, we approximate $\log(1+z)$ around $z=0$ as $\log(1+z)=z-z^{2}/2+z^{3}/3-\cdots$. Therefore, we have